Shift Summary for eve shift, 31/Jan/00


Subject: Shift Summary for eve shift, 31/Jan/00
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:48:25 CST


              Shift Summary for eve shift, 31/Jan/00

On Shift:PSh, JAWS (with RK,DJ,RR,AG,ST)

Hours of beam available at all intensities:5.7 hrs
Hours of beam available at nominal intensity:5.7 hrs
Hours of KAMI data (special runs not included):0.5 hr
Hours of special runs (laser scans, muon runs...):4.3 hrs

Accounting of lost time: TeV cryo problems. Run initialization problems.

========================================================================
Runs ended this shift:
Run 15719 - KAMI_HADRON (for target scan)
Run 15723 - KAMI_MINBIAS (in progress)

Tapes brought to FCC this shift:
  (including tapes for online split)

========================================================================
Accesses during shift:
AG/JW to turn on AGEL HV.
RF/PSh to NM2 to hunt for beam loss.

========================================================================
Overview of shift: Finished target scan study. Start long min-bias run.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
                     15:30 Take over day shift target scan studies
                                   just as beam disappears. Cryo problems.
                                   RickC puts NM21V,NM2HV back to nominal.

                     16:00 SashaG/JW make access to turn on AGEL HV.

                     16:40 RickF/PSh make access to NM2 to look
                                   for evidence of beam loss. No source
                                   found.

                     18:20 Beam returns. Run initialization problems.
                                   Dart fresh start.

                     19:00 Start run 15719. Finish beam target scan.
                                   Beam returned to nominal.
                                   Results of horizontal targetting scans
                                   shown in KAMI log I (p.200) No major
                                   shape change seen, possibly due to steep
                                   targetting angle.

                     19:43 Start run 15723 - KAMI_MINBIAS run with
                                   no absorber.
                                      

========================================================================
Follow up Notes on Previous Shift

Finished target scan.

========================================================================
Items to Follow up on Next Shift

Continue KAMI_MINBIAS running.



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:48:29 CST